OpenStudy (anonymous):

the position vector of a particle is r(t) find the requested vector the velocity at t=2 for r(t)=(2t^2+5t+6)i-3t^3j+(10-t^2)k

OpenStudy (agreene):

remember that r'(t)=v(t) thus: d/dt [r(t)]=v(t) => d/dt <2t^2+5t+6,-3t^3,10-t^2> \[<\frac{d}{dt}2t^2+5t+6,\frac{d}{dt}-3t^3,\frac{d}{dt}10-t^2>\] \[r'(t)=v(t)=<4t+5,-9t^2,-2t>\] or, in vector hat notation: \[v(t)=4t+5 i-9t^2j-2tk\] now you have velocity with respect to time, simply plug and chug for the value.

OpenStudy (anonymous):

Thank u